A 10-gram bullet fired at 100 m/s collides with a block of wood at rest. The mass of the block is 490 grams. The collision is inelastic. What is the speed of the block and bullet after collision? (Unit in m/s)
1
Expert's answer
2020-11-25T07:14:58-0500

Inelastic collision implies that after the interaction, the object previously separated move as one. So, applying conservation of momentum, we have:


"mv+MV=(M+m)u,\\\space\\
u=\frac{mv+MV}{M+m}=\frac{0.01\cdot100+0.49\cdot0}{0.49+0.01}=2\text{ m/s}."